### v1

# Three quantities arising from Bézout’s identity and resultants of integer polynomials

Abstract.
In this paper, we study three quantities arising naturally from Bézout’s identity, the resultant and the reduced resultant of two non-zero coprime integer polynomials. We establish several new divisibility relations among them. We also pose two conjectures by making computations.
Key words and phrases:
Integer polynomial, Bézout’s identity, resultant, reduced resultant
2020 Mathematics Subject Classification:
11C08, 11R09
1. Introduction
In this paper, integer polynomials are meant to be polynomials in $\mathbb{Z}[x]$, that is, polynomials in the variable $x$ and with integer coefficients. They are definitely among the fundamental objects in algebra and number theory. The research around their arithmetic properties has a very long history and is still very active. The arithmetic properties include reducibility, degeneracy, decomposability, moduli of roots, signature of roots, multiplicative dependence of roots, etc. See [1, 3, 4, 5, 6, 9, 10, 11, 12, 13, 14, 15, 20] for some recent developments.
In this paper, we study the relations among three quantities arising naturally from Bézout’s identity, the resultant and the reduced resultant of two non-zero coprime integer polynomials. Here, we say two integer polynomials are coprime if they are coprime in the polynomial ring $\mathbb{Q}[x]$ over the rational numbers $\mathbb{Q}$.
From now on, $f(x)$ and $g(x)$ are two non-zero coprime integer polynomials. Below we define the three quantities studied in this paper with respect to $f(x)$ and $g(x)$.
By Bézout’s identity, we know that there exist uniquely two polynomials $u(x)$ and $v(x)$ in $\mathbb{Q}[x]$ such that:
(1.1) $u(x)f(x)+v(x)g(x)=1,\quad\deg(u)<\deg(g),\quad\deg(v)<\deg(f).$
Then, we define the quantity $B(f,g)$ to be the least common multiple of the denominators of all the coefficients of $u(x)$ and $v(x)$ in (1.1) (the coefficients are reduced to their lowest terms). Letting $B=B(f,g)$, we have
(1.2) $Bu(x)f(x)+Bv(x)g(x)=B,\quad Bu(x),Bv(x)\in\mathbb{Z}[x].$
Since $f(x)$ and $g(x)$ are non-zero and coprime, their resultant $\mathrm{Res}(f,g)$ is a non-zero integer. Then, we define $R(f,g)=|\mathrm{Res}(f,g)|$. It is well known that the resultant $\mathrm{Res}(f,g)$ can be expressed as a linear combination of $f(x)$ and $g(x)$ over the polynomial ring $\mathbb{Z}[x]$; see [7, page 157, Proposition 9]. So, there are two integer polynomials $s(x),t(x)$ such that
(1.3) $s(x)f(x)+t(x)g(x)=R(f,g).$
In addition, we define the quantity $r(f,g)$ to be the smallest positive integer in the following set
$f(x)\mathbb{Z}[x]+g(x)\mathbb{Z}[x]=\left\{u(x)f(x)+v(x)g(x):\,u(x),v(x)\in\mathbb{Z}[x]\right\}.$
This quantity is called reduced resultant in [22]. In addition, it is also called congruence number in [23, Definition 2.6], and moreover, the authors in [23] also described its applications.
By definition, $r(f,g)$ is the smallest positive integer in the intersection $(f(x)\mathbb{Z}[x]+g(x)\mathbb{Z}[x])\cap\mathbb{Z}$, which in fact is an ideal of the integer ring $\mathbb{Z}$ and so is equal to $r(f,g)\mathbb{Z}$. Thus, by definition we directly have
(1.4) $r(f,g)\mid B(f,g)\quad\textrm{and}\quad r(f,g)\mid R(f,g).$
This is known to the experts.
Moreover, for the ideal generated by $f(x)$ and $g(x)$ in $\mathbb{Z}[x]$, if we compute its Gröbner basis (see [7]), then the constant polynomial in this basis is exactly $r(f,g)$. In this paper, we compute $r(f,g)$ based on this fact and using Mathematica.
From the discussions in MathOverflow (see, for instance, [19, 24]), one can see that many people are interested with the relations among the three quantities $B(f,g)$, $r(f,g)$ and $R(f,g)$. However, so far we haven’t found a paper studying their relations systematically.
In this paper, we want to investigate systematically the divisibility relations among the three quantities $B(f,g)$, $r(f,g)$ and $R(f,g)$. We believe that some of them are new; such as Theorems 3.1, 4.1, 5.1 and 5.6. We also pose two conjectures at the end.
2. Preliminaries
In this section, we present some notation and some results which are needed later on.
First, we recall that for an integer polynomial $h(x)$, the content of $h$, denoted by $\mathrm{cont}(h)$, is the greatest common divisor of the coefficients of $h$. Besides, we denote by $L(h)$ the leading coefficient of $h$.
For two integer polynomials $h(x)$ and $q(x)$, recall that their resultant $\mathrm{Res}(h,q)$ is defined to be the determinant of their Sylvester matrix. We first list three classical properties of resultants as follows (see [7] or [17] for their proofs):
(2.1) $\mathrm{Res}(q,h)=(-1)^{\deg(h)\deg(q)}\mathrm{Res}(h,q),$
(2.2) $\mathrm{Res}(h,q_{1}q_{2})=\mathrm{Res}(h,q_{1})\mathrm{Res}(h,q_{2}),$
(2.3) $\mathrm{Res}(c,h)=\mathrm{Res}(h,c)=c^{\deg(h)},$
where $h,q,q_{1},q_{2}$ are integer polynomials, and $c$ is a non-zero constant. The result in [8, Lemma 4.1] asserts that if we can write $h=sq+t$ with integer polynomials $s,t$, then
(2.4) $\mathrm{Res}(q,h)=L(q)^{\deg(h)-\deg(t)}\mathrm{Res}(q,t).$
For two non-zero coprime integer polynmials $f(x),g(x)$, we have known that $R(f,g)$ can be expressed as a linear combination of $f$ and $g$ in $\mathbb{Z}[x]$ (see (1.3)). The following result gives a more restricted expression of $R(f,g)$.
Theorem 2.1.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d$ be the greatest common divisor of their leading coefficients. Then, there are two unique polynomials $u(x)$ and $v(x)$ such that $\deg(u)<\deg(g)$, $\deg(v)<\deg(f)$, $d\mid\mathrm{cont}(u)$, $d\mid\mathrm{cont}(v)$, and
$u(x)f(x)+v(x)g(x)=R(f,g).$
In particular, $d\mid R(f,g)$.
Proof.
First, we write
$f(x)=a_{n}x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0},$
$g(x)=b_{m}x^{m}+b_{m-1}x^{m-1}+\cdots+b_{1}x+b_{0},$
with $a_{n}b_{m}\neq 0$. Regarding the Sylvester matrix of $f(x)$ and $g(x)$, for each $i$ ($11\leq i\leq n+m-$ 1), multiply the $i$-th column by $x^{n+m-i}$ and add it to the $(n+m)$-th column (here we view $x$ as an arbitrary non-zero integer). Then, we obtain
$\mathrm{Res}(f,g)=\det\left(\begin{array}[]{ccccccc}a_{n}&a_{n-1}&a_{n-2}&\cdots&0&0&x^{m-1}f\\ 0&a_{n}&a_{n-1}&\cdots&0&0&x^{m-2}f\\ 0&0&a_{n}&\cdots&0&0&x^{m-3}f\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&a_{1}&a_{0}&xf\\ 0&0&0&\cdots&a_{2}&a_{1}&f\\ b_{m}&b_{m-1}&b_{m-2}&\cdots&0&0&x^{n-1}g\\ 0&b_{m}&b_{m-1}&\cdots&0&0&x^{n-2}g\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&b_{1}&b_{0}&xg\\ 0&0&0&\cdots&b_{2}&b_{1}&g\end{array}\right).$
By the definition of $d$, we can write $a_{n}=ad,b_{m}=bd$ for two coprime positive integers $a,b$. Then, we obtain
$\mathrm{Res}(f,g)$ $=d\cdot\det\left(\begin{array}[]{ccccc}a&a_{n-1}&\cdots&0&x^{m-1}f\\ 0&a_{n}&\cdots&0&x^{m-2}f\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&a_{1}&f\\ b&b_{m-1}&\cdots&0&x^{n-1}g\\ 0&b_{m}&\cdots&0&x^{n-2}g\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&b_{1}&g\end{array}\right).$
We denote by $A$ the determinant on the right hand side of the above identity. For $1\leq i,j\leq m+n$, we denote the $(i,j)$-th cofactor of $A$ by $A_{ij}$. Then, expanding $A$ via cofactors along the last column, we get
$\mathrm{Res}(f,g)$ $=d\Big{(}x^{m-1}fA_{1,m+n}+\cdots+fA_{m,m+n}+$
$\qquad\qquad x^{n-1}gA_{m+1,m+n}+\cdots+gA_{m+n,m+n}\Big{)}$
$=d(x^{m-1}A_{1,m+n}+\cdots+A_{m,m+n})f+$
$\qquad\qquad d(x^{n-1}A_{m+1,m+n}+\cdots+A_{m+n,m+n})g$
$=u_{1}(x)f(x)+v_{1}(x)g(x),$
where
$u_{1}(x)=d(x^{m-1}A_{1,m+n}+...+A_{m,m+n})\in\mathbb{Z}[x],$
$v_{1}(x)=d(x^{n-1}A_{m+1,m+n}+\cdots+A_{m+n,m+n})\in\mathbb{Z}[x].$
Taking absolute value, we obtain
(2.5) $R(f,g)=|\mathrm{Res}(f,g)|=u(x)f(x)+v(x)g(x),$
where $u(x)=u_{1}(x)$ and $v(x)=v_{1}(x)$ if $\mathrm{Res}(f,g)>0$, and otherwise $u(x)=-u_{1}(x)$ and $v(x)=-v_{1}(x)$.
Note that in the above deductions, we view $x$ as an arbitrary integer. That is, (2.5) holds for any integer $x$. So, (2.5) also holds if we view $x$ as a variable.
Noticing that each cofactor $A_{i,m+n}$ ($1\leq i\leq m+n$) is an integer, clearly we have
$d\mid\mathrm{cont}(u),\qquad d\mid\mathrm{cont}(v).$
Meanwhile, it can be seen that $\deg(u)\leq m-1<\deg(g)$, $\deg(v)\leq n-1<\deg(f)$. Finally, since $f$ and $g$ are coprime, one can prove by contradiction that $u(x)$ and $v(x)$ are unique. ∎
3. Relations between $B(f,g)$ and $r(f,g)$
In this section, we investigate the relation between $B(f,g)$ and $r(f,g)$ for two non-zero coprime integer polynomials $f(x),g(x)$. We have already known that $r(f,g)\mid B(f,g)$ by (1.4). The following result is about the converse of this divisibility relation.
Theorem 3.1.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, there exists a non-negative integer $k$ such that
$B(f,g)\mid d^{k}r(f,g),$
where $d$ is the greatest common divisor of the leading coefficients of $f(x)$ and $g(x)$.
Proof.
For simplicity, let $B=B(f,g)$ and $r=r(f,g)$. By definition, there exist two integer polynomials $u(x)$ and $v(x)$ such that
(3.1) $u(x)f(x)+v(x)g(x)=r.$
Since $r\in\mathbb{Z}$, we must have
(3.2) $\deg(u)+\deg(f)=\deg(v)+\deg(g).$
So, either $\deg(u)<\deg(g)$ and $\deg(v)<\deg(f)$, or $\deg(u)\geq\deg(g)$ and $\deg(v)\geq\deg(f)$.
If $\deg(u)<\deg(g)$ and $\deg(v)<\deg(f)$, then we have
$\frac{u(x)}{r}f(x)+\frac{v(x)}{r}g(x)=1.$
According to the definition of $B$ (that is, $B(f,g)$), we have $B\mid r$.
Now, we assume that $\deg(u)\geq\deg(g)$ and $\deg(v)\geq\deg(f)$. Recall that $L(f)$ and $L(g)$ are the leading coefficients of $f$ and $g$ respectively. Let $d=\gcd(L(f),L(g))$. From (3.1) and noticing $r\in\mathbb{Z}$, we have
$L(u)\cdot L(f)=-L(v)\cdot L(g),$
where $L(u)$ and $L(v)$ are the leading coefficients of $u$ and $v$ respectively. Moreover, we get
(3.3) $L(u)\cdot\frac{L(f)}{d}=-L(v)\cdot\frac{L(g)}{d}.$
Since $\gcd(\frac{L(f)}{d},\frac{L(g)}{d})=1$, we have $\frac{L(f)}{d}\mid L(v)$ and $\frac{L(g)}{d}\mid L(u)$. Then, we define two integer polynomials:
$h_{1}(x)=\frac{dL(u)}{L(g)}x^{\deg(u)-\deg(g)},\qquad q_{1}(x)=\frac{dL(v)}{L(f)}x^{\deg(v)-\deg(f)}.$
By (3.2) and (3.3), we obtain
(3.4) $h_{1}(x)+q_{1}(x)=0.$
So, we have
$d\cdot u(x)=g(x)h_{1}(x)+u_{1}(x),$
$d\cdot v(x)=f(x)q_{1}(x)+v_{1}(x),$
where $u_{1}(x)$ and $v_{1}(x)$ are some integer polynomials such that $\deg(u_{1})<\deg(u)$ and $\deg(v_{1})<\deg(v)$. Combining the above two formulas with (3.1), we get
$f(x)g(x)(h_{1}(x)+q_{1}(x))+u_{1}(x)f(x)+v_{1}(x)g(x)=dr,$
which, together with (3.4), gives
$u_{1}(x)f(x)+v_{1}(x)g(x)=dr.$
Repeating the above steps, after a finite number of steps (say $k$ steps), there must exist integer polynomials $u_{k}(x)$ and $v_{k}(x)$ such that $\deg(u_{k})<\deg(g)$ and $\deg(v_{k})<\deg(f)$, and
$u_{k}(x)f(x)+v_{k}(x)g(x)=d^{k}r.$
Hence, as before we obtain
$B\mid d^{k}r.$
This completes the proof. ∎
We remark that the non-negative integer $k$ in Theorem 3.1 is at most $\max(\deg(u)-\deg(g)+1,0)$ with polynomial $u$ in (3.1). Moreover, one can find an upper bound for $\deg(u)$ in [2, Theorem A].
When the leading coefficients of $f(x)$ and $g(x)$ are coprime (that is, $d=1$ in Theorem 3.1), we have $B(f,g)\mid r(f,g)$. Then, noticing $r(f,g)\mid B(f,g)$, we directly get the following corollary.
Corollary 3.2.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, if their leading coefficients are coprime, then $B(f,g)=r(f,g)$.
If the leading coefficients of $f(x)$ and $g(x)$ are not coprime, then either $B(f,g)=r(f,g)$ or $B(f,g)\neq r(f,g)$. Both cases can happen; see Examples 3.3 and 3.4.
Example 3.3.
Let $f(x)=6x^{3}-6x^{2}-6x-6$ and $g(x)=6x^{3}-6x^{2}-6x+5$. Then, $B(f,g)=11$, $R(f,g)=287496$ and $r(f,g)=11$, and so, $R(f,g)=6^{3}r(f,g)^{3}$.
Example 3.4.
Let $f(x)=6x^{2}+5$ and $g(x)=6x^{2}-4x+1$. Then, $B(f,g)=22=2\times 11$, $R(f,g)=1056=2^{5}\times 3\times 11$ and $r(f,g)=11$.
We also remark that if the leading coefficients of $f(x)$ and $g(x)$ are not coprime and let $d$ be their greatest common divisor, then $B(f,g)$ and $r(f,g)$ may not have the same prime factors, and also $B(f,g)$ and $dr(f,g)$ may not have the same prime factors; see Example 3.4. However, by Theorem 3.1 we know that any prime factor of $B(f,g)$ divides $dr(f,g)$.
4. Relations between $r(f,g)$ and $R(f,g)$
For any two non-zero coprime integer polynomials $f(x),g(x)$, we have already known that $r(f,g)\mid R(f,g)$ by (1.4). The following result is about the converse of this divisibility relation.
Theorem 4.1.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d$ be the greatest common divisor of their leading coefficients. Then, we have
$R(f,g)\mid d^{k}r(f,g)^{\max(n,m)},$
where $n=\deg(f),m=\deg(g),k=\deg(uf)=\deg(vg)$ with polynomials $u,v$ in (3.1).
Proof.
Let $R=R(f,g)$ and $r=r(f,g)$. Since $r$ is a non-zero integer, by (2.3) we have
$\mathrm{Res}(f,r)=r^{\deg(f)}=r^{n}.$
In addition, by (3.1), there exist two integer polynomials $u,v$ such that
$uf+vg=r,$
which, together with (2.2) and (2.4), gives
$\begin{split}\mathrm{Res}(f,r)&=\mathrm{Res}(f,uf+vg)=L(f)^{0-k}\mathrm{Res}(f,vg)\\ &=L(f)^{-k}\mathrm{Res}(f,g)\mathrm{Res}(f,v),\end{split}$
where $k=\deg(vg)=\deg(uf)$. So, we have
$\mathrm{Res}(f,r)=r^{n}=L(f)^{-k}\mathrm{Res}(f,g)\mathrm{Res}(f,v),$
which implies
(4.1) $L(f)^{k}r^{n}=\mathrm{Res}(f,g)\mathrm{Res}(f,v).$
Similarly, we obtain
$\mathrm{Res}(g,r)=r^{m}=L(g)^{-k}\mathrm{Res}(g,f)\mathrm{Res}(g,u),$
which implies
(4.2) $L(g)^{k}r^{m}=\mathrm{Res}(g,f)\mathrm{Res}(g,u).$
Note that by definition and by (2.1), we have
$R=|\mathrm{Res}(f,g)|=|\mathrm{Res}(g,f)|.$
By the definition of $d$, we write $L(f)=ad,L(g)=bd$ for two coprime positive integers $a,b$ (that is $\gcd(a,b)=1$). Then, combining (4.1) with (4.2), we get
(4.3) $R\mid a^{k}d^{k}r^{n},\qquad R\mid b^{k}d^{k}r^{m}.$
Now, for any prime factor $p$ of $R$, if $p^{j}\mid R$ and $p^{j+1}\nmid R$ for some positive integer $j$, then by (4.3) and noticing $\gcd(a,b)=1$, we must have $p^{j}\mid d^{k}r^{\max(n,m)}$. Hence, we have
$R\mid d^{k}r^{\max(n,m)}.$
This completes the proof. ∎
If $f(x)$ is monic (that is $L(f)=1$), by (4.1) we recover a special case of a result in [21]:
Corollary 4.2.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$ with $f(x)$ monic, we have $R(f,g)\mid r(f,g)^{n}.$
By Theorem 4.1 and noticing $r(f,g)\mid R(f,g)$ and $d\mid R(f,g)$ (see Theorem 2.1), we directly get that $R(f,g)$ and $dr(f,g)$ have the same prime factors.
Corollary 4.3.
For two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d=\gcd(L(f),L(g))$. Then, $R(f,g)$ and $dr(f,g)$ have the same prime factors.
We remark that when $d>1$, it can happen that $R(f,g)$ and $r(f,g)$ do not have the same prime factors; see Example 3.3.
5. Relations between $B(f,g)$ and $R(f,g)$
Given two non-zero coprime integer polynomials $f(x)$ and $g(x)$, it is known to the experts that $B(f,g)\mid R(f,g)$. We enhance this divisibility relation in the following result.
Theorem 5.1.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d$ be the greatest common divisor of their leading coefficients. Then, we have
$dB(f,g)\,\Big{|}\,R(f,g).$
Proof.
Let $B=B(f,g)$ and $R=R(f,g)$. First, from (1.2) we know that there exist uniquely two integer polynomials $u_{1}(x)$ and $v_{1}(x)$ such that $\deg(u_{1})<\deg(g)$, $\deg(v_{1})<\deg(f)$ and
(5.1) $u_{1}(x)f(x)+v_{1}(x)g(x)=B.$
By the definition of $B$, we have
(5.2) $\gcd(\mathrm{cont}(u_{1}),\mathrm{cont}(v_{1}))=1.$
In addition, by Theorem 2.1 there exist two integer polynomials $u_{2}(x)$ and $v_{2}(x)$ such that $\deg(u_{2})<\deg(g)$, $\deg(v_{2})<\deg(f)$ and:
(5.3) $u_{2}(x)f(x)+v_{2}(x)g(x)=\frac{R}{d}.$
Here, we let $R_{0}=R/d$.
Now, from (5.1) and (5.3), we obtain
$\frac{u_{1}(x)}{B}f(x)+\frac{v_{1}(x)}{B}g(x)=1,$
$\frac{u_{2}(x)}{R_{0}}f(x)+\frac{v_{2}(x)}{R_{0}}g(x)=1.$
By the uniqueness of the polynomials $u(x),v(x)$ in (1.1), we must have
$\frac{u_{1}(x)}{B}=\frac{u_{2}(x)}{R_{0}},\qquad\frac{v_{1}(x)}{B}=\frac{v_{2}(x)}{R_{0}}.$
So, we obtain
$B\mathrm{cont}(u_{2})=R_{0}\mathrm{cont}(u_{1}),\qquad B\mathrm{cont}(v_{2})=R_{0}\mathrm{cont}(v_{1}).$
If $B\nmid R_{0}$, then there is a prime number $p$ such that $p\mid B$, $p\mid\mathrm{cont}(u_{1})$ and $p\mid\mathrm{cont}(v_{1})$, and so this contradicts with (5.2). Hence, we must have $B\mid R_{0}$. This completes the proof. ∎
When $d>1$, by Theorem 5.1 we directly get $B(f,g)\neq R(f,g)$, and then also $r(f,g)\neq R(f,g)$.
Corollary 5.2.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, if their leading coefficients are not coprime, then $B(f,g)\neq R(f,g)$ and $r(f,g)\neq R(f,g)$.
If the leading coefficients of $f(x)$ and $g(x)$ are coprime, then either $B(f,g)=R(f,g)$ or $B(f,g)\neq R(f,g)$. Both cases can happen; see the following two examples.
Example 5.3.
Let $f(x)=2x^{3}-x^{2}-x$ and $g(x)=x^{3}-x^{2}+x+1$. Then, $B(f,g)=r(f,g)=R(f,g)=2$.
Example 5.4.
Let $f(x)=2x^{3}+x^{2}-x-1$ and $g(x)=x^{3}-x^{2}+x+1$. Then, $B(f,g)=r(f,g)=3$ and $R(f,g)=27$.
Moreover, combining Corollary 3.2 with Corollary 5.2, we directly obtain the following result.
Corollary 5.5.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, then $B(f,g)=R(f,g)$ if and only if $r(f,g)=R(f,g)$.
Now, we establish a converse of the divisibility relation in Theorem 5.1 as follows.
Theorem 5.6.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d$ be the greatest common divisor of their leading coefficients. Then, we have
$R(f,g)\mid d^{n+m-1}B(f,g)^{\max(n,m)},$
where $n=\deg(f)$ and $m=\deg(g)$.
Proof.
Let $B=B(f,g)$ and $R=R(f,g)$. We have known that there exist uniquely two integer polynomials $u_{1}(x)$ and $v_{1}(x)$ such that $\deg(u_{1})<\deg(g)$, $\deg(v_{1})<\deg(f)$ and
$u_{1}(x)f(x)+v_{1}(x)g(x)=B.$
Then, applying the same arguments as in the proof of Theorem 4.1, we obtain
$R\mid d^{k}B^{\max(n,m)},$
where $k=\deg(u_{1}f)=\deg(v_{1}g)\leq n+m-1$. This completes the proof. ∎
By Theorems 5.1 and 5.6, we directly get that $R(f,g)$ and $dB(f,g)$ have the same prime factors.
Corollary 5.7.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d=\gcd(L(f),L(g))$. Then, $R(f,g)$ and $dB(f,g)$ have the same prime factors.
We remark that when $d>1$, it can happen that $B(f,g)$ and $R(f,g)$ do not have the same prime factors; see Example 3.4.
Combining Corollary 4.3 with Corollary 5.7, we directly get the following corollary.
Corollary 5.8.
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, if their leading coefficients are coprime, then $B(f,g)$, $r(f,g)$ and $R(f,g)$ have the same prime factors.
6. Numerical data
In this section, we present some numerical data to see how often we have $B(f,g)=r(f,g)$ (or moreover $R(f,g)=r(f,g)$) for two coprime integer polynomials $f(x),g(x)$ which are given randomly.
First, for an integer polynomial $h(x)\in\mathbb{Z}[x]$, we denote by $\mathrm{H}(h)$ the maximum of the absolute values of its coefficients. Recall that $L(h)$ stands for the leading cofficient of $h$.
Then, for three positive integers $n,m$ and $H$, we define the following set of pairs of coprime integer polynomials:
$S_{n,m}(H)=$$\Big{\{}(f,g)\in\mathbb{Z}[x]^{2}:\,\deg f=n,\deg g=m,\gcd(f,g)=1,$
$\quad L(f)>0,L(g)>0,\mathrm{H}(f)\leq H,\mathrm{H}(g)\leq H\Big{\}}.$
Note that in the set $S_{n,m}(H)$, both $f$ and $g$ are assumed to have positive leading coefficients, because
$f(x)\mathbb{Z}[x]+g(x)\mathbb{Z}[x]=(\pm 1)f(x)\mathbb{Z}[x]+(\pm 1)g(x)\mathbb{Z}[x].$
In Table 6.1, each value corresponds to $n,m,H$ is the following percentage:
|$\frac{|\{(f,g)\in S_{n,m}(H):\,B(f,g)=r(f,g)\}|}{|S_{n,m}(H)|}\times 100\%,$
which are rounded to the nearest two decimal places in the table. Recall that the probability that two randomly given integers are coprime is $6/\pi^{2}\approx 0.6079$; see [18, Theorem 332]. So, by Corollary 3.2, we can roughly say that the above percentage is at least $60\%$ when $H$ is sufficiently large. However, Table 6.1 shows that this percentage is much larger (more than $90\%$). Moreover, when $(n,m)=(1,1)$, this percentages corresponding to $H=30,50$ and $100$ are respectively: $90.29\%$, $90.32\%$, and $90.11\%$.
Hence, we pose the following conjecture. It suggests that for two coprime integer polynomials $f(x),g(x)$ given randomly, we usually have $B(f,g)=r(f,g)$.
Conjecture 6.1.
For any positive integers $n$ and $m$, we have
$\lim_{H\to\infty}\frac{|\{(f,g)\in S_{n,m}(H):\,B(f,g)=r(f,g)\}|}{|S_{n,m}(H)|}\geq\frac{9}{10}.$
$(n,m)$ $H$ 2 3 4 5 6
(1, 1) 97.44% 95.90% 93.61% 94.73% 92.13%
(1, 2) 98.58% 98.04% 96.53% 97.48% 95.98%
(1, 3) 99.36% 99.09% 98.13% 98.81% 98.04%
(2, 2) 99.00% 99.10% 98.08% 98.78% 98.09%
(2, 3) 99.52% 99.55% 99.04% 99.43% 99.03%
(3, 3) 99.69% 99.78% 99.51% 99.72% 99.51%
Table 6.1. Percentage of $B(f,g)=r(f,g)$
From Corollary 5.5, we know that the frequency of $B(f,g)=R(f,g)$ is the same as that of $r(f,g)=R(f,g)$. So, we only need to consider how often we have $B(f,g)=R(f,g)$.
In Table 6.2, each value corresponds to $n,m,H$ is the following percentage:
|$\frac{|\{(f,g)\in S_{n,m}(H):\,B(f,g)=R(f,g)\}|}{|S_{n,m}(H)|}\times 100\%,$
which are rounded to the nearest two decimal places in the table. As the above, by Corollary 5.2 we can roughly say that this percentage is at most $61\%$ when $H$ is sufficiently large. However, Table 6.2 shows that this percentage is much smaller except the case when $(n,m)=(1,1)$. Table 6.2 suggests that for two coprime integer polynomials $f(x),g(x)$ given randomly, we often have $B(f,g)\neq R(f,g)$ when $\deg f$ and $\deg g$ are large.
Regarding Table 6.2, we pose the following conjecture.
Conjecture 6.2.
For any positive integers $n$ and $m$ with $(n,m)\neq(1,1)$, we have
$\lim_{H\to\infty}\frac{|\{(f,g)\in S_{n,m}(H):\,B(f,g)\neq R(f,g)\}|}{|S_{n,m}(H)|}>\frac{1}{2}.$
Given two coprime integer polynomials of degree one $f(x)=ax+b$ and $g(x)=cx+d$, we directly have $R(f,g)=|ad-bc|$ and
$\frac{c}{bc-ad}\cdot f(x)-\frac{a}{bc-ad}\cdot g(x)=1.$
So, by definition we have $B(f,g)=|ad-bc|/\gcd(a,c)$. Hence, $B(f,g)=R(f,g)$ whenever $\gcd(a,c)=1$. This can explain why the percentages in Table 6.2 corresponding to the case when $(n,m)=(1,1)$ are large.
$(n,m)$ $H$ 2 3 4 5 6
(1, 1) 82.05% 81.54% 75.31% 79.47% 70.35%
(1, 2) 64.15% 61.14% 51.73% 58.16% 46.13%
(1, 3) 62.80% 60.44% 50.56% 57.52% 45.04%
(2, 2) 46.14% 49.81% 41.47% 49.08% 37.76%
(2, 3) 46.61% 49.54% 41.33% 48.66% 37.66%
(3, 3) 46.29% 49.24% 41.24% 48.25% 37.74%
Table 6.2. Percentage of $B(f,g)=R(f,g)$
Acknowledgement
For the research, Z. Liu and X. Li were supported by the Guangdong College Students’ Innovation and Entrepreneurship Training Program (No. S202410574067); M. Sha was supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515010635).